OpenStudy (anonymous):

Verify the Identity cot(x + pi/2) = -tan x

OpenStudy (anonymous):

@AriPotta @iGreen Please help :/

OpenStudy (anonymous):

I know that pi/2 would be 0 and the equation would end up like this cot x = -tan x but I do not know what to do after that...

OpenStudy (anonymous):

@satellite73 Could you help me out after you're done with that question?

OpenStudy (freckles):

do you know cotangent function or even the tangent function or both odd functions?

OpenStudy (anonymous):

cotangent function as in cot x= 1/tan x? and tan x = 1/ cot x?

OpenStudy (freckles):

ok odd function as in tan(-x)=-tan(x) and so on.. but nevermind you could write cot in terms of cos and sin and use the sum identities

OpenStudy (freckles):

\[\frac{\cos(x+\frac{\pi}{2})}{\sin(x+\frac{\pi}{2})}=?\]

OpenStudy (freckles):

expand the top and bottom

OpenStudy (freckles):

i will brb i think @TylerD is here to help also

OpenStudy (tylerd):